Find the remaining parts of a triangle with β=50^0 , b=5, and c=6

Answer by stanbon(75887)

Find the remaining parts of a triangle with β=50^0 , b=5, and c=6
---
Using the Law of Sines, solve for angle C:
---
sin(C)/6 = sin(50)/5
---
sin(C) = (6/5)sin(50)
--
sin(C) = 0.9193
---
C = arcsin(0.9193)
---
C = 66.82 degrees
----------------
Then A = 1-(50+66.82)
A = 63.18 degrees
--------------------
Now, using Law of Sines, solve for
"a" and for "b".
====================
Cheers,
Stan H.
====================